Institut for Matematik og Datalogi Odense Universitet March 2, 1995

JFB

DM15 – F95 – Week 20

Lecture, February 27

We finished chapter 9 and through section 10.2 of chapter 10 in Sudkamp.

Lecture, March 6

We will finish chapter 10, section 6.9, and the first two sections of chapter 11 in Sudkamp.

Lecture, March 13

We will continue in chapter 11 in Sudkamp, possibly finishing the chapter.

Problems to be discussed in week 11

• Do problem 7 on page 249 of Sudkamp.

• Do problem 8 on page 249 of Sudkamp. Note that homomorphisms are defined in problem 9 on page 182.

• Do problem 15 on page 250 of Sudkamp.

• Do problem 2 on page 277 of Sudkamp.

• Do problem 4 on page 277 of Sudkamp.

• Do problem 6 on page 277 of Sudkamp.